Faz, Inc., manufactures and sells two products: Product X0 and Product W7. Data concerning the expected production of each product and the expected total direct labor-hours (DLHs) required to produce that output appear below:

Expected Production Direct Labor-Hours Per Unit Total Direct Labor-Hours
Product X0 2,000 4 8,000
Product W7 440 3 1,320
Total direct labor-hours 9,320

The direct labor rate is $23.60 per DLH. The direct materials cost per unit is $148.50 for Product X0 and $130 for Product W7.

The company is considering adopting an activity-based costing system with the following activity cost pools, activity measures, and expected activity:

Estimated Expected Activity
Activity Cost Pools Activity Measures Overhead Cost Product X0 Product W7 Total
Labor-related DLHs $ 276,078 8,000 1,320 9,320
Production orders orders 19,348 540 740 1,280
Order size MHs 241,614 4,040 4,140 8,180
$ 537,040

Which of the following statements concerning the unit product cost of Product W7 is true? (Round your intermediate calculations to 2 decimal places.)

rev: 03_25_2018_QC_CS-119201

Multiple Choice

  • The unit product cost of Product W7 under traditional costing is greater than its unit product cost under activity-based costing by $219.37.

  • The unit product cost of Product W7 under traditional costing is less than its unit product cost under activity-based costing by $41.39.

  • The unit product cost of Product W7 under traditional costing is greater than its unit product cost under activity-based costing by $41.39.

  • The unit product cost of Product W7 under traditional costing is less than its unit product cost under activity-based costing by $219.37.
Answer #1
Traditional costing:
Estimated overhead cost 537040
Divide by Total Direct Labor-Hours 9320
Traditional costing overhead rate 57.62
Product W7 Overhead per unit 172.86 =57.62*3
Activity based costing:
Activity Cost Pools Activity rate
Labor-related 29.62 =276078/9320
Production orders 15.12 =19348/1280
Order size 29.54 =241614/8180
Activity Cost Pools Product W7
Labor-related 39098.4 =1320*29.62
Production orders 11188.8 =740*15.12
Order size 122295.6 =4140*29.54
Total overhead 172582.8
Divide by Units 440
Product W7 Overhead per unit 392.23
Difference under both methods 219.37 =392.23-172.86
The unit product cost of Product W7 under traditional costing is less than its unit product cost under activity-based costing by $219.37.
Option D is correct